pint dunsy nch ge by e)i 04 ind E (x-): unuihy chon grven by elsewhace
Answer #1

1)

E(X-Y)=\int _0^1\int _0^1\left(x-y\right)\left(x+y\right)dxdy=0

\int _0^1\left(x-y\right)\left(x+y\right)dx=-y^2+\frac{1}{3}

\int _0^1\left(-y^2+\frac{1}{3}\right)dy=0

2)

P(X > 0.1 , Y > 0.1) =\int _{0.1}^1\int _{0.1}^1x+ydxdy=0.891

\int _{0.1}^1x+ydx=0.9y+0.495

\int _{0.1}^1\left(0.9y+0.495\right)dy=0.891

Please rate